An Affordable Particle Detector For Education

Brendan James Arnold November 12, 2004

Abstract

This project continues ongoing work done at Bristol University researching an affordable, safe particle detector to be used as experimental apparatus in ’A Level’ education. In particular we built and tested Resistive Plate Chambers (RPCs). Typical configurations used in high energy physics use a gas mixture that includes freon and isobutane. We obtained results avoiding the use of freon which is a greenhouse gas and also looked at possible replacements for isobutane which is highly flammable.

Introduction

All main UK A-Level syllabus(Ref) feature particle physics at some point. Some feature cosmic rays in particular (Ref). Commercially available particle detectors seem to be restricted to cloud chambers and although there are relatively cheap (Ref) ones available, they need a source of dry ice, a resource that many schools do not have. A further problem with cloud chambers is that data is difficult to log. Atlhough a video camera arrangement can be used, this is relatively costly in terms of equipment and requires time-consuming human analysis to get results. This also closes off the possibility of more exciting experiments such as the Swedish SEASA schools collaboration where particle detectors from different schools are linked into a huge detector array (Ref).

We considered many types of detector to investigate. Work has already been done on scintillation based detectors and is being continued in a parallel project by a colleague at Bristol (Ref) however the latest estimates for costs are still quite high, XXX for a cloud chamber and XXX for scintillation based detectors (Ref). We decided to investigate an entirely new avenue for the Bristol project and look into Resistive Plate Chambers (RPCs) chiefly because of the apparent low cost of materials for the chamber and the relatively simple design principle.

Most research into RPCs is for use in high energy physics laboratories and other industrial environments and not the classroom where the equipment will be in close proximity to schoolchildren. There are many inherant safety risks involved in RPCs which we aim to minimise in this project. In addition we have to take into account that, should the project be successful, our design will be produced in large numbers, we should try and avoid the use of harmful greenhouse gases which are prevailent in most current industrial RPCs.

A Basic RPC Design

An RPC is closely related to the spark chamber in that it registers charge freed from ionising particles. The key difference is that a resistive plate does not spark due to the highly resistive plates of glass in between the anode and cathode. This has the advantage that lower voltages can be used. A particle event is registered by ’pick-up’ strips of metal which lie outside the electrodes.

Figure (Ref) illustrates the basic RPC principle. A passing ionising particle ionises the gas locally along its path in the gas gap. The resulting electrons and ions move towards opposite electrodes and are ’slowed’ by the two inner resistive layers. Since directly above and below the particle’s path there is briefly a concentration of either positive or negative charge, in accordance to the laws of electrostatics an opposite charge is induced in the

Figure 1: Schematics of a basic RPC

readout strips. Note that the readout strips are isolated from the electrodes by another resistive layer.

Originally the resistive layers were bakelite which had the right resistivity properties but had to be treated with linseed oil to obtain a smooth plane surface(Ref). Smoother surfaces help reduce electric noise [1] More recently float glass has been used since this has similar resistivity properties and does not need treating. It is also inexpensive and readily available in large sheets. The electrodes are typically graphite and applied to the outer surfaces of the resistive layers as paint. Graphite electrodes which have a fairly high resistance in comparison to, say, silver also prevent the dispersion of the charge from happening too rapidly to reigister.

The gas mixture varies greatly between RPC designs, however they all have the following key properties,

1. Must be ionised only by particles of high enough energy as to be cosmic rays. Since cosmic ray energies are usually very high indeed this role is typically fulfilled by a noble gas such as Argon.

2. Must absorb residual ultraviolet light from the ionisation process to prevent further ionisation elswhere in the chamber. This gas is known as the ‘quencher’.

3. Must have a high electron affinity to prevent too much lateral spread of electrons in the chamber.

Much research has gone into the gas mixture. Typical quenchers are Freon, Sulpher Hexaflouride, Tetraflouroethane, IsobutaneOxygen and and progress has been made towards finding a mix that is not highly flammable and does not include greeehouse gases.

Bibliography

[1] Study of an Avalanche-mode Resistive Plate Chamber, J Ying et al. J. Phys. G: Nucl. Part. Phys. 26 (2000) page 1292